OpenStudy (brittany):

solver for the variable. Reduce the answer to lowest terms. 5/6m=-30

OpenStudy (anonymous):

1/m=-30*6/5=-36 1+36m=0 m=-1/36

OpenStudy (anonymous):

i take it this is \[\frac{5}{6}m=-30\] multiply both sides by the reciprocal of \[\frac{5}{6}\] which is \[\frac{6}{5}\] to get \[\frac{6}{5}\times \frac{5}{6}m=-30\times \frac{6}{5}\] \[m=\frac{-30\times 6}{5}=\frac{-6\times 6}{5}=-\frac{36}{5}\]